# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ IX Destruction of static-order phase via droplets
In finite systems, a ferromagnetic-like state with static order is observed both in 2D and 3D. As system size increases, however, our numerical simulations indicate that this static state is destabilized. This is supported by Figs. 10 and 13 (right parts) showing that the $\tilde{J}$-transition-point between static order and oscillations in finite systems (indicated by a sharp change in the slope of $R$ and $\mathcal{L}$ becoming zero) increases with system size in a way that does not seem to converge to a finite $\tilde{J}$. Figures 14 and 20a showing $R$ and $\mathcal{L}$ as a function of $\tilde{K}$ have a very small region of static order, when $\tilde{K}$ is close to zero. However, this region too shrinks when $L$ increases.
@@
The static-order phase is destabilized by nucleation of droplets that grow and flip the magnetization, in alternating order of A- and B-spins. This mechanism is demonstrated in Fig. 23, where snapshots of the spins at different times are shown for increasing system sizes, in a 2D system. The full evolution is shown in Movie 4. For $L=10$ (Fig. 23a), both species do not change their magnetization. There are attempts of B-spins to nucleate droplets, but those shrink and disappear. For $L=40$ (Fig. 23b), the system is no longer static: a handful of B-spins droplets (one clearly observed in the figure) surpass a certain critical droplet size, grow and flip the magnetization. Then, the same occurs for the A-spins, and so on, leading to an oscillatory behavior of $M_{A}(t)$ and $M_{B}(t)$, but with intermediate periods in which $M_{A}(t)$ and $M_{B}(t)$ are constants. When $L=200$ (Fig. 23c), multiple droplets ($\gg 1$) drive the magnetization flipping of each species, making the $M_{A}(t)$ and $M_{B}(t)$ curves smooth at all times, i.e, there is no finite period of time in which the magnetization is constant and the system could be called static. We note that the figure shows a 2D system but a similar behavior is observed in 3D.
@@
Fig. 23: Instability of the static order phase in the thermodynamic limit due to droplet nucleation. Top four rows: Snapshots of $\sigma^{A}$ and $\sigma^{B}$ in a 2D system with linear system size (a) $L=10$, (b) $L=40$, (c) $L=200$. Time evolves from top to bottom. Fifth row: the corresponding total magnetizations, $M_{A}$ and $M_{B}$, as a function of time. Obtained from MC simulation with parameters $\tilde{J}=2.8$ and $\tilde{K}=0.3$, and with initialization at $t=0$ with all spins up.
@@
To see why non-reciprocity allows droplets to grow, in contrast with the equilibrium Ising model where they tend to shrink, assume first that the system is in the static-order phase, so most spins in both lattices are up (Fig. 24a) [92]. Since the system is static, it can be mapped into two equilibrium Ising models with opposite magnetic fields, with energy
@@
where $H\approx K$ for A-spins and $H\approx-K$ for B-spins (see Eq. (11) and Fig. 24b). While sub-system A is in a global free energy minimum in the effective equilibrium system, sub-system B is in a metastable state, as it prefers a state with opposite magnetization. Sub-system $B$ then transitions to its global minimum by nucleating droplets larger than a critical size $\rho_{c}$ above which the effective magnetic field, which acts on the volume of the droplet, overcomes the nearest neighbor attraction, imposing surface tension, which makes the droplet grow [196, 92] (Fig. 24c). After a droplet has expanded beyond $\rho_{c}$, the stable sub-system becomes metastable and nucleates droplets, and so on (Fig. 24d). Crucially, due to the symmetry between $A$ and $B$ spins, both droplet types expand with the same radial velocity. Therefore, a nested droplet remains much smaller than its “mother-droplet” until the latter finishes expanding (see Sec. XII in which this symmetry is absent). A simplified view of this mechanism can be found in defining a “selfish free energy” for the zero-dimensional MF equation, see Appendix B.
Fig. 24: Schematic drawing of the droplet argument for the instability of the static-order phase. (a) A system in the static-order phase can be reduced into (b) two equilibrium Ising models subject to magnetic fields, where one system is in a stable equilibrium while the other in a metastable state. (c) The metastable system nucleates droplets that expand and flip the magnetization. (d) The stable system becomes metastable and nucleates droplets that expand with the same velocity, and so on.
The argument sketched above is contingent upon $\rho_{c}$ being finite, which is the case in any finite dimension (and finite couplings $\tilde{J}$ and $\tilde{K}$). Hence, we conjecture that in the fully anti-symmetric NR Ising model, the static-order phase is unstable in any finite dimension.
IX.1 Between static-order and swap: the droplet regime
@@
In this regime, as $\tilde{J}$ increases (and as $\tilde{K}$ decreases), the critical droplet size increases, and as a results the droplets become more sparse and the distance they travel before colliding with each other increases. Eventually, $\tilde{J}$ becomes large enough (or $\tilde{K}$ small enough) that the critical droplet size becomes larger than the system size. In that case, the magnetization can flip when fluctuations are of the order of the system size, in the same way as in the usual equilibrium Ising model.
@@
This is demonstrated in Fig. 25 where 2D-slice snapshots of the $\theta$ field, as well as the magnetizations of the two species as a function of time, are shown for a 3D system with linear size $L=80$ at increasing values of $\tilde{J}$. For small $\tilde{J}$, the 3D system is in a noisy homogeneous swap phase. There are no apparent droplets but rather noisy patterns without a well-defined shape (Fig. 25a,e). Put differently, independent fluctuations do not have time to expand significantly before colliding. For higher $\tilde{J}$ values we observe droplets that drive the magnetization flips (Fig. 25b-c,f-g). For an even larger $\tilde{J}$, no droplets are observed and the magnetizations do not change over a long period of time (Fig. 25d,h).
@@
Fig. 25: Transition from swap to static order in a finite 3D system, as a function of $\tilde{J}$. Top: 2D-slice simulation snapshots of the $\theta$ field for varying values of $\tilde{J}$. Bottom: The corresponding total magnetizations $M_{A}$ (blue) and $M_{B}$ (red) as a function of time measured in simulation sweeps. (a,e) $\tilde{J}=1.6$. (b,f) $\tilde{J}=2.2$. (c,g) $\tilde{J}=2.3$. (d,h) $\tilde{J}=2.5$. Other parameters are $\tilde{K}=0.3$ and $L=80$. As $\tilde{J}$ increases, a noisy homogeneous swap phase is replaced by droplet-induced swap and then becomes static and ordered.
@@
The droplet regime is characterized by a distinct behavior of the oscillation period as a function of system size. In Fig. 26a, the oscillation period in 3D, $T_{\rm osc}$, is shown as a function of $\tilde{J}$ and a fixed $\tilde{K}$, for increasing $L$ values. In the non-droplet regime, i.e. small $\tilde{J}$, $T_{\rm osc}$ is independent of system size while in the droplet-induced swap regime (large $\tilde{J}$) $T_{\rm osc}$ decreases with system size at intermediate $L$, until it converges in the $L\to\infty$ limit to a finite value, independent of $L$ (see overlap of orange and green lines in Fig. 26a up to $\tilde{J}\approx 2.3$). In the large $\tilde{J}$ limit, and for fixed system size, droplets are so rare that each magnetization flip is driven by a single droplet. In this regime, the droplet can emerge in any lattice position at a given time step, thus $T_{\rm osc}$ scales with $L^{-3}$ [196, 92] (see Fig. 26b).
@@
The different oscillation regimes share similarities with the relaxation of the magnetization in an Ising model from an initial state magnetized opposite to the applied field $H$. Reference [196] identifies four regimes, depending on the field strength and system size. The “strong-field” regime in which “the droplet picture is inappropriate”, and three other regimes which are classified using droplet theory by the relation between three length scales: the critical droplet radius $\rho_{c}$, the mean droplet separation $\rho_{0}$, and the linear system size $L$. These are the multi-droplet regime $L\gg\rho_{0}\gg\rho_{c}$, the single-droplet regime $\rho_{0}\gg L\gg\rho_{c}$, and the “coexistence region” $\rho_{0}\gg\rho_{c}\gg L$ where the system behaves as if there is no applied field. The average metastable lifetime calculated in Ref. [196] is shown to be independent of system size in strong field and multi-droplet regime yet highly dependent on system size in the single-droplet and coexistence regime, in a way reminiscent of Fig. 26a. We note that the analogy drawn here is not perfect because non-reciprocal interactions are different from an externally applied field in that they change once the magnetization flips.
@@
Fig. 26: Behavior of the oscillation period in the swap phase. (a) The oscillation period of the total magnetizations $T_{\rm osc}$ in the 3D system, on a logarithmic scale, as a function of $\tilde{J}$ for fixed $\tilde{K}=0.3$ and different system sizes. (b) The same with a rescaled oscillation period, $L^{3}T_{\rm osc}$. Dashed grey line in the top panel separates a regime in which the oscillations are driven by well-defined droplets (right) from a regime in which they do not (left). Obtained from MC simulations.
@@ IX.2 The fate of the droplet regime
In 3D, the scenario is more complex. While droplets persist at large system size, distant regions become less synchronized and the system loses its global order, see Fig. 27. While $R$ clearly decreases (see right region in Fig. 13), it does not converge with system size. Thus, it is unclear to what extent order is lost in the thermodynamic limit, and whether it remains finite or is ultimately deemed disordered ($R=0$). If the latter is correct, it remains an open question what is the type of phase transition that occurs and where lies the critical line between swap and disorder in this droplet regime. We remark that the decrease in order depicted in Fig. 27, suggests a coexistence of domains with different $\theta$ values (where $\theta$ is uniform within each domain), that may indicate a first-order-like transition. However, the domains are not static: they change over time due to droplet nucleation, growth, and collisions.
Fig. 27: Loss of synchronization in the droplet regime with increased system size. Snapshots of the $\theta$ field in 2D slices taken from a 3D MC simulation with $L=80,160$, and $320$. Other system parameters are $\tilde{K}=0.3$ and $\tilde{J}=2.25$. As system size increases, more $\theta$ values coexistence (seen as more colors in a single snapshot).
The decrease in order in the droplet-induced-swap regime causes an intriguing behavior, observed only for large system size, ($L\geq 160$): synchronization is non-monotonic as a function of $\tilde{J}$ (or $\tilde{K}$). See $L=320$ line in Fig. 13, where $R$ increases monotonically beyond the Hopf bifurcation until at $\tilde{J}\sim 2.05$, and then starts decreasing. It then reaches a local minimum at $\tilde{J}\sim 2.25$ and increases again until saturation at $\tilde{J}\sim 2.55$. A similar behavior is observed in the left part Fig. 20a, for $L=320$, where $R$ is non-monotonic as a function of $\tilde{K}$. In both cases the anomalous decrease starts when the swap phase transitions from noisy homogeneous swap to droplet-induced swap.
The mechanism under which global synchronization decreases with increased $\tilde{J}$ (or decreased $\tilde{K}$) in the droplet regime is shown in Fig. 28, which compares two simulations with the same $\tilde{K}=0.3$ value: one at the onset of droplet-induced swap ($\tilde{J}=2.04$), and the other well within the droplet-induced swap ($\tilde{J}=2.24$). Simulation snapshots (Figure 28a) reveal that while the system is more correlated over short distances at larger $\tilde{J}$ values, it loses global order due to large regions that are out-of-phase with the rest of the system (Fig. 28a). This can be quantified by a size dependent synchronization parameter,
@@
where the sum is over all sites inside a box of linear size $l\leq L$, and the average is taken with respect to different boxes and time. The size dependent synchronization parameter is shown in Fig. 28b, for squares instead of cubes to match with the 2D snapshots. For small $l$, $R(l)$ is larger for the higher $\tilde{J}$ value while for large $l$, $R(l)$ is larger for the smaller $\tilde{J}$ value. Moreover, while $R(l)$ at low $\tilde{J}$ decays exponentially with $l$, in the high $\tilde{J}$ regime it does not and the second derivative of $R(l)$ changes signs. This means that within the droplet-induced swap regime, the system cannot be described by a single correlation length, and instead has multiple length scales related to the droplet parameters such as the critical droplet size, the typical distance between droplets, and the correlation length inside a single droplet.
Fig. 28: Decrease in global order in the droplet regime as a function of $\tilde{J}$. (a) Snapshots of 2D slices taken from a 3D simulation with $\tilde{K}=0.3$ and $L=400$, comparing two different $\tilde{J}$ values, $\tilde{J}=2.04$ (left) and $\tilde{J}=2.24$ (right). The top snapshots are taken shortly after initialization while the bottom snapshots are taken after 10000 MC sweeps. (b) The synchronization parameter $R(l)$ defined in Eq. (62), measured at steady state as a function of $l$, for the two $\tilde{J}$ values in panel a. Averages are taken over $(L/l)^{2}$ squares and 1000 MC sweeps.

### v2

@@ IX Destruction of the static-order phase via droplets
In finite systems, a ferromagnetic-like state with static order is observed both in 2D and 3D. As system size increases, however, our numerical simulations indicate that this static state is destabilized. This is supported by Figs. 9e-f and 12a-b (right parts) showing $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ (for a fixed $\tilde{K}$). The $\tilde{J}$-transition-point between static order and oscillations in finite systems (indicated by a sharp change in the slope of $R$ and $\mathcal{L}$ becoming zero) increases with system size in a way that does not seem to converge to a finite $\tilde{J}$. Figures 12c and 18a showing $R$ and $\mathcal{L}$ as a function of $\tilde{K}$ (for a fixed $\tilde{J}$) have a very small region of static order (when $\tilde{K}$ is close to zero) and it is hard to see what is its fate in the thermodynamic limit. However, this region seems to shrink when $L$ increases in Fig 18a, suggesting a destabilization of the static-order phase.
@@
The static-order phase is destabilized by nucleation of droplets that grow and flip the magnetization, in alternating order of A- and B-spins. This mechanism is demonstrated in Fig. 21, where snapshots of the spins at different times are shown for increasing system sizes, in a 2D system. The full evolution is shown in Movie 4. For $L=10$ (Fig. 21a), both species do not flip their magnetization. There are attempts of B-spins to nucleate droplets, but those shrink and disappear. For $L=40$ (Fig. 21b), the system is no longer static: a handful of B-spins droplets (one clearly observed in the figure) surpass a certain critical droplet size, grow and flip the magnetization. Then, the same occurs for the A-spins, and so on, leading to an oscillatory behavior of $M_{A}(t)$ and $M_{B}(t)$, but with intermediate periods in which $M_{A}(t)$ and $M_{B}(t)$ are constants. When $L=200$ (Fig. 21c), multiple droplets ($\gg 1$) drive the magnetization flipping of each species, making the $M_{A}(t)$ and $M_{B}(t)$ curves smooth at all times, i.e, there is no finite period of time in which the magnetization is constant and the system could be called static. We note that the figure shows a 2D system but a similar behavior is observed in 3D.
@@
Fig. 21: Instability of the static-order phase in the thermodynamic limit due to droplet nucleation. Top four rows: Snapshots of $\sigma^{A}$ and $\sigma^{B}$ in a 2D system with linear system size (a) $L=10$, (b) $L=40$, (c) $L=200$. Time evolves from top to bottom. See Movie 4 for the full evolution. Fifth row: the corresponding total magnetizations, $M_{A}$ and $M_{B}$, as a function of time. Obtained from MC simulations with parameters $\tilde{J}=2.8$ and $\tilde{K}=0.3$, and with initialization at $t=0$ with all spins up.
@@
To see why nonreciprocity allows droplets to grow, in contrast with the equilibrium Ising model where they tend to shrink, assume first that the system is in the static-order phase, so most spins in both lattices are up (Fig. 22a) [96]. We further assume $K$ is very small, which is when static order is most likely to be stable. As long as the system is static, it can be mapped into two equilibrium Ising models with opposite magnetic fields, with energy
@@
where $H\approx K$ for A-spins and $H\approx-K$ for B-spins (see Eq. (11) and Fig. 22b). While sub-system A is in a global free energy minimum in the effective equilibrium system, sub-system B is in a metastable state, as it prefers a state with opposite magnetization. Provided that the system size is large enough, sub-system $B$ then transitions to its global minimum by nucleating droplets larger than a critical size $\rho_{c}$ above which the effective magnetic field, which acts on the volume of the droplet, overcomes the nearest neighbor attraction, imposing surface tension, which makes the droplet grow [233, 96, 99, 234] (Fig. 22c). After a droplet has expanded beyond $\rho_{c}$, the stable sub-system becomes metastable and nucleates droplets, and so on (Fig. 22d). Crucially, due to the symmetry between $A$ and $B$ spins, both droplet types expand with the same speed. Therefore, a nested droplet remains much smaller than its “mother-droplet” until the latter reaches the system size or collides with other droplets (see Sec. XII in which this symmetry is absent). A simplified view of this mechanism can be found in defining a “selfish free energy” for the zero-dimensional mean-field equation, see Appendix E.
Fig. 22: Droplet argument for the instability of the static-order phase. (a) A system in the static-order phase can be reduced into (b) two equilibrium Ising models subject to magnetic fields, where one system is in a stable equilibrium while the other in a metastable state. (c) The metastable system nucleates droplets that expand and flip the magnetization. (d) The stable system becomes metastable and nucleates droplets that expand with the same velocity, and so on.
The argument sketched above relies on $\rho_{c}$ being finite, which is the case in any finite dimension (and finite couplings $\tilde{J}$ and $\tilde{K}$) because the volume term ultimately dominates at sufficiently large droplet sizes. Hence, we conjecture that in the fully anti-symmetric nonreciprocal Ising model, the static-order phase is unstable in the thermodynamic limit in any finite dimension.
IX.1 Between static order and swap: the droplet regime
@@
In the droplet-induced swap regime, the critical droplet size increases as $\tilde{J}$ increases (and as $\tilde{K}$ decreases). As a result, the droplets become more sparse and the distance they travel before colliding with each other increases. Eventually, $\tilde{J}$ becomes large enough (or $\tilde{K}$ small enough) that the critical droplet size becomes larger than the system size. In that case, the magnetization can flip when fluctuations are of the order of the system size, in the same way as in the usual equilibrium Ising model.
@@
This is demonstrated in Fig. 23 where 2D-slice snapshots of the $\theta$ field, as well as the magnetizations of the two species as a function of time, are shown for a 3D system with linear size $L=80$ at increasing values of $\tilde{J}$ (and decreasing values of $\tilde{K}$). For small $\tilde{J}$ (high $\tilde{K}$), the 3D system is in a noisy homogeneous swap phase. There are no apparent droplets but rather noisy patterns without a well-defined shape (Fig. 23a,e). Put differently, independent fluctuations do not have time to expand significantly before colliding. For larger $\tilde{J}$ (smaller $\tilde{K}$) values we observe droplets that drive the magnetization flips (Fig. 23b-c,f-g). For an even larger $\tilde{J}$ (smaller $\tilde{K}$), no droplets are observed and the magnetizations do not change over a long period of time (Fig. 23d,h).
@@
Fig. 23: Transition from swap to static order in a finite 3D system, as a function of $\tilde{J}$ and $\tilde{K}$. (a-d) Top: 2D-slice snapshots of the $\theta$ field, obtained from MC simulations with increasing values of $\tilde{J}$. Bottom: The corresponding total magnetizations $M_{A}$ (blue) and $M_{B}$ (red) as a function of time measured in simulation sweeps. (a) $\tilde{J}=1.6$. (b) $\tilde{J}=2.2$. (c) $\tilde{J}=2.3$. (d) $\tilde{J}=2.5$, all with $\tilde{K}=0.3$ and $L=80$. As $\tilde{J}$ increases, a noisy homogeneous swap phase is replaced by droplet-induced swap and then becomes static and ordered. (e-f) The same as panels a-d (top) but with $\tilde{J}=1.6$ and decreasing values of $\tilde{K}$. (e) $\tilde{K}=0.3$. (f) $\tilde{K}=0.08$. (g) $\tilde{K}=0.05$. (h) $\tilde{K}=0.02$.
@@
The droplet regime is characterized by a distinct behavior of the oscillation period, $T_{\rm osc}$, as a function of system size. In Fig. 24a, the oscillation period in 3D is shown as a function of $\tilde{J}$ with a fixed $\tilde{K}$, for increasing $L$ values. In the non-droplet regime, i.e. small $\tilde{J}$, $T_{\rm osc}$ is independent of system size (note that the period is finite at the critical point $\tilde{J}_{c}$, supporting a Hopf-like behavior), while in the droplet-induced swap regime (large $\tilde{J}$) $T_{\rm osc}$ decreases with system size at intermediate $L$ but converges in the $L\to\infty$ limit to a finite value, independent of $L$ (see overlap of orange and green lines in Fig. 24a up to $\tilde{J}\approx 2.3$). At larger $\tilde{J}$ values, in a finite system, droplets are so rare that each magnetization flip is driven by a single droplet (here we assume that $\tilde{J}$ is not too large, such that the critical droplet size is still smaller than the system size). In this regime, the single droplet can emerge in any lattice position at a given time step, thus $T_{\rm osc}$ scales with $L^{-3}$ [233, 96], see Fig. 24b.
@@
The different oscillation regimes share similarities with the relaxation of the magnetization in an Ising model from an initial state magnetized opposite to the applied field $H$. Reference [233] identifies four regimes, depending on the field strength and system size. The “strong-field” regime in which “the droplet picture is inappropriate”, and three other regimes which are classified using droplet theory by the relation between three length scales: the critical droplet radius $\rho_{c}$, the mean droplet separation $\rho_{0}$, and the linear system size $L$. These are the multi-droplet regime $L\gg\rho_{0}\gg\rho_{c}$, the single-droplet regime $\rho_{0}\gg L\gg\rho_{c}$, and the “coexistence region” $\rho_{0}\gg\rho_{c}\gg L$ where the system behaves as if there is no applied field. The average metastable lifetime calculated in Ref. [233] is shown to be independent of system size in strong field and multi-droplet regime yet highly dependent on system size in the single-droplet and coexistence regime, in a way reminiscent of Fig. 24a. We note that the analogy drawn here is not perfect because nonreciprocal interactions are different from an externally applied field in that they change once the magnetization flips.
@@
Fig. 24: Behavior of the oscillation period in the 3D swap phase. (a) The oscillation period of the total magnetizations $T_{\rm osc}$, on a logarithmic scale, as a function of $\tilde{J}$ for fixed $\tilde{K}=0.3$ and different system sizes. (b) The same with a rescaled oscillation period, $L^{3}T_{\rm osc}$. Dashed grey line in panel a separates a regime in which the oscillations are driven by well-defined droplets (right) from a regime in which they do not (left). $\tilde{J}_{c}$ marked on the x-axis is the critical point for the disorder-to-swap phase transition. The period is finite at $\tilde{J}_{c}$, supporting a Hopf-like behavior. Obtained from MC simulations.
@@ IX.2 The fate of the droplet regime
In 3D, the scenario is more complex. While droplets persist at large system size, distant regions become less synchronized and the system loses its global order. This is illustrated in Fig. 25 which shows snapshots of the $\theta$ field, taken from simulations in the droplet-induced swap regime, with different system sizes. In Fig. 25, we see that at system sizes $L=(80,160,320)$ there are regions with $n=(2,3,4)$ distinct $\theta$-valued (i.e. $n$ different colors in the plot), respectively. This decreases the synchronization. Indeed, the synchronization order parameter, $R$, is shown to decrease with $L$, in the right part of Fig. 12a, corresponding to the droplet-induced swap region. While $R$ decreases with system size, it does not converge. Thus, it is unclear to what extent order is lost in the thermodynamic limit, and whether it remains finite or is ultimately deemed disordered ($R=0$). If the latter is correct, it remains an open question what is the type of phase transition that occurs and where lies the critical line between swap and disorder in this droplet regime.
Fig. 25: Loss of synchronization in the droplet regime with increased system size. Snapshots of the $\theta$ field (color code in Fig. 2) in 2D slices taken from a 3D MC simulation with $L=80,160$, and $320$. Other system parameters are $\tilde{K}=0.3$ and $\tilde{J}=2.25$. We see $n=(2,3,4)$ distinct $\theta$-valued regions for $L=(80,160,320)$, respectively.
The decrease in order in the droplet-induced-swap regime causes an intriguing behavior, observed only for large system size, ($L\geq 160$): synchronization is non-monotonic as a function of $\tilde{J}$ (or $\tilde{K}$). See $L=320$ line in Fig. 12a, where $R$ increases monotonically with $\tilde{J}$ beyond the Hopf bifurcation until $\tilde{J}\sim 2.05$, and then starts decreasing. It then reaches a local minimum at $\tilde{J}\sim 2.25$ and increases again until saturation at $\tilde{J}\sim 2.55$. A similar behavior is observed in the left part Fig. 18a, for $L=320$, where $R$ is non-monotonic as a function of $\tilde{K}$. In both cases the anomalous decrease starts when the swap phase transitions from noisy homogeneous swap to droplet-induced swap.
The mechanism under which global synchronization decreases with increased $\tilde{J}$ (or decreased $\tilde{K}$) in the droplet regime is shown in Fig. 26, which compares two simulations with the same $\tilde{K}=0.3$ value: one at the onset of droplet-induced swap ($\tilde{J}=2.04$), and the other well within the droplet-induced swap ($\tilde{J}=2.24$). Simulation snapshots (Figure 26a) reveal that while the system is more correlated over short distances at larger $\tilde{J}$ values, it loses global order due to large regions that are out-of-phase with the rest of the system. This can be quantified by a size-dependent synchronization parameter,
@@
where the sum is over all sites inside a box of linear size $l\leq L$, and the average is taken with respect to different boxes and time. The size dependent synchronization parameter is shown in Fig. 26b, for squares instead of cubes to match with the 2D snapshots. For small $l$, $R(l)$ is larger for the system with the higher $\tilde{J}$, whereas for large $l$, $R(l)$ is larger for the system with the lower $\tilde{J}$. Moreover, while $R(l)$ at low $\tilde{J}$ decays exponentially with $l$, in the high $\tilde{J}$ regime it does not and the second derivative of $R(l)$ changes signs. This means that within the droplet-induced swap regime, the system cannot be described by a single correlation length, and instead has multiple length scales related to the droplet parameters such as the critical droplet size, the typical distance between droplets, and the correlation length inside a single droplet.
Fig. 26: Decrease in global order in the droplet regime as a function of $\tilde{J}$. (a) Snapshots of 2D slices taken from a 3D simulation with $\tilde{K}=0.3$ and $L=400$, comparing two different $\tilde{J}$ values, $\tilde{J}=2.04$ (left) and $\tilde{J}=2.24$ (right). The top snapshots are taken shortly after initialization while the bottom snapshots are taken after 10000 MC sweeps. (b) The synchronization parameter $R(l)$ defined in Eq. (71), measured in steady state as a function of $l$, for the two $\tilde{J}$ values in panel a. Averages are taken over $(L/l)^{2}$ squares and 1000 MC sweeps.